# Exploratory Notes on Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

### v3

@@ Abstract
I present a coordinate-free, symbolic framework for deciding whether a given set of polygonal faces can form a closed, genus-zero polyhedral surface and for predicting how such a surface could be decomposed into internal tetrahedra. The method uses only discrete incidence variables, such as the number of internal tetrahedra $T$, internal gluing triangles $N_{i}$, and internal triangulation segments $S_{i}$, and applies combinatorial feasibility checks before any geometric embedding is attempted.
For polyhedra in normal form, I record exact incidence identities linking $V,E,F$ to a flatness parameter $S:=\sum_{f}(\operatorname{deg}f-3)$, and I identify parity-sensitive effects in $E$, $F$, and $S$. The external identities and parity-sensitive bounds hold universally for genus-0 polyhedral graphs. For internal quantities, I prove exact relations $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (with $S_{i}$ taken to be the number of interior edges) and obtain restricted linear ranges within a shell–aligned ladder subclass (SALT), where at most one interior edge is introduced per layer. Consequently, I propose a symbolic workflow that yields rapid pre-checks for structural impossibility, reducing the need for costly geometric validation in computational geometry, graphics, and automated modeling.
@@ 1 Introduction and Background
Scope and fixed assumptions: Throughout this work, I limit attention to genus-zero (simply connected) surfaces, with no geometric embeddability tests (i.e., no attempt to prove realizability in $\mathbb{R}^{3}$), and without consideration of higher-genus or non-manifold topologies. Unless stated otherwise, all internal “range” results are given under the SALT+MIE restriction (Definition 2.1), i.e. normal form with at most one new interior segment per layer, introduced only when necessary. All identities and bounds presented are symbolic rather than metric, and some are conjectural or empirically validated rather than proved in full generality.
@@
Combining these with the constraint that all vertices have degree at least 33 yields sharp bounds:
@@
The auxiliary variables NN, $M$, $S$, $N_{i}$, and $S_{i}$ do not appear in the classical Euler–Steinitz framework. They are introduced here to support symbolic decomposition bookkeeping and to connect surface composition to volumetric tetrahedralization. N represents the total internal angle units on the external surface, with one unit defined as $\pi$ radians ($180^{\circ}$ ). $M$ is derived from polygon types and measures “external angle fullness” in the same units. $S$ counts the non-intersecting triangulation segments added to polygonal faces to produce a fully triangulated surface and also serves as a discrete flatness measure. $N_{i}$ counts the number of internal gluing angle units, each corresponding to a shared triangular face between two tetrahedra. $S_{i}$ counts the internal non-intersecting triangulation segments required to complete a volumetric mesh during decomposition. None of these auxiliary quantities are edges in the polyhedron’s graph; they exist only in the symbolic model and may not correspond to realizable geometric features.
@@ 1.1 Illustrative Examples of $N_{i}$, $S$, and $S_{i}$
Figure 1: Rectangular pyramid illustrating the flatness measure $S$. The single internal diagonal on the quadrilateral base (orange) is the unique surface triangulation segment needed to obtain a fully triangulated boundary, hence $S=1$. This corresponds to one coplanarity merge relative to the maximally triangulated case and, by Proposition 3.3, decreases both $F$ and $E$ by 11 while preserving $V$ and $\chi=2$.
@@ 2.6 Tetrahedral decomposition and trivalent vertex prioritization
Definition 2.1 (Shell–aligned ladder tetrahedralizations (SALT)).
In essence, I restrict attention to genus–0, normal–form decompositions built by peeling a single outer shell of tetrahedra at a time (a layer). Each layer is created by coning existing boundary triangles and, only if necessary, introducing at most one new interior segment between boundary vertices to complete gluing across that layer. No interior vertices (Steiner points) and no edge bisections are allowed. Internal adjacencies occur only along triangular faces.
Definition 2.2 (Minimal Interior Edge (MIE) rule).
A new interior segment is introduced only when strictly necessary to complete the current layer; otherwise none is added.
@@ 3.1 Vertex Basis and Initial Angle Unit Calculation
For a genus-0 polyhedron in normal form, consider first the triangulated case ($S=0$). In this case every face has degree 33, so the incidence relation is
@@
Now, increasing the flatness parameter $S$ by one corresponds to replacing a triangular face by a face of degree 44; this removes exactly one edge and one face from the triangulation. More generally, each unit of $S$ reduces both $E$ and $F$ by exactly one compared to the triangulated case. Therefore
@@
Let $G$ be the 11-skeleton of a genus–0 polyhedron (simple, planar, and 33-connected) with $V\geq 4$ vertices and $E$ edges. Then
@@
    $V$ is even and all vertices are trivalent (degree 33); or
@@
    $V$ is odd and all but one vertex are trivalent, the remaining vertex being 44-valent.
@@
Every vertex has degree at least 33, so by the handshake lemma
@@
If $V$ is even and all vertices have degree 33, then $2E=3V$ and $E=3V/2$, attaining the bound.
If $V$ is odd and all but one vertex are trivalent and the remaining vertex is 44-valent, then 2E=3(V1)+4=3V+12E=3(V-1)+4=3V+1, so $E=(3V+1)/2$, again attaining the bound.
@@
Moreover, as $S$ increases by 11 (flattening step), both $E$ and $F$ decrease by 11, and for $k=0,1,\dots,S_{\max}$ one has E=3V6k,F=2V4k.E=3V-6-k,\;F=2V-4-k.
@@
Start from any maximally planar (triangulated) 33–connected planar graph on $V$ vertices (e.g., a stacked triangulation), which has $S=0$, E=3V6E=3V-6, F=2V4F=2V-4. A “coplanarity merge” of two adjacent boundary triangles corresponds combinatorially to deleting their shared edge, which decreases both $E$ and $F$ by 11 and keeps $V$ fixed. In a stacked triangulation one can choose a set of V4V-4 edges (e.g., along successive stacked faces) whose deletion preserves simplicity, planarity, and 33–connectivity; performing $S$ such merges yields a graph with the claimed $(E,F)$. ∎
@@ 3.3 Configuration Series over Flatness
From Proposition 3.1, each unit decrease in the flatness measure $S$ increases both $F$ and $E$ by 11. Starting from the extremal values in Corollary 3.2, I have:
@@ 3.4 Counting Distinct Combinatorial Configurations
Heuristic 3.1 (External bounds and parity effects).
@@
since each tetrahedron contributes four triangle units and every internal gluing triangle removes two boundary triangle units by pairing. These two expressions for NN are consistent with Proposition 5.4 and Proposition 3.1: indeed $F+S=4T-2N_{i}$. This connection between external angle counts and internal tetrahedral composition will be used in §5 to link internal volumetric parameters directly to external face and edge counts.
@@ 3.5 Triangulation Sequences and the Invariance of Euler’s Characteristic
If $S$ denotes the number of coplanarity merges of adjacent triangles (each merge decreases both $F$ and $E$ by 11 while $V$ is fixed), then for all $S$ in the admissible range,
@@
Each triangle has three edges and each edge is shared by two triangles, so $2E_{\max}=3F_{\max}$. Substituting $E_{\max}=\tfrac{3}{2}F_{\max}$ into Euler’s formula $V-E_{\max}+F_{\max}=2$ yields $F_{\max}=2(V-2)$ and $E_{\max}=3(V-2)$. A coplanarity merge of two adjacent triangles removes one surface edge and one triangle, so $F\mapsto F-1$ and $E\mapsto E-1$ with $V$ unchanged. Hence VE+FV-E+F is invariant and equal to 22 for all admissible $S$. ∎
@@ 4.4 Count of Valid Internal Configurations
Heuristic 4.1 (Internal ranges and configuration ladder).
@@ 5.1 Exact topological Euler link
  • Faces: $N+N_{i}$, where NN is the number of boundary triangular faces (angle units),
@@ 5.3 Heuristic extension of Euler’s characteristic
Heuristic 5.1 (Extended Euler-type identity).
@@ 6 A Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility
If NN is odd, then $V$ is non-integer and no valid genus-0 polyhedron in normal form with such a face–degree distribution can exist.
@@ 10 Acknowledgements
This appendix records (i) proved external and topological identities used in the main text, and (ii) the empirical/heuristic identities and bounds collected from worked decompositions. All variables are symbolic and coordinate-free. Genus 0 and normal-form assumptions apply where stated. All linear “max” values quoted below are valid within the SALT+MIE class (Definition 2.1). Non–SALT triangulations, such as dense pulling triangulations of cyclic polytopes, can exceed these bounds by creating quadratically many tetrahedra without adding interior vertices.
@@ E.1 Proved External Identities and Bounds
(Equation (E.6) counts triangle units: each tetrahedron contributes 44, each internal gluing triangle cancels 22.)
@@ E.2 Proved Topological Identity for 3D Complexes
(Prop. 5.1). In particular, when $V_{i}=0$ (normal form, no interior vertices),

### v5

# Symbolic Constraints in Polyhedral Enclosure and Tetrahedral Decomposition in Genus-0 Polyhedra

@@ Abstract
I present a coordinate-free, symbolic framework for determining whether a given set of polygonal faces can form a closed, genus-zero polyhedral surface and for predicting how such a surface could be decomposed into internal tetrahedra. The method uses only discrete incidence variables, such as the number of internal tetrahedra $T$, internal gluing triangles $N_{i}$, and internal triangulation segments $S_{i}$, and applies combinatorial feasibility checks before any geometric embedding is attempted.
For polyhedra in normal form, I record exact incidence identities linking $V,E,F$ to a flatness parameter $S:=\sum_{f}(\operatorname{deg}f-3)$, and I identify parity-sensitive effects in $E$, $F$, and $S$. The external identities and parity-sensitive bounds hold universally for genus-0 polyhedral graphs. For internal quantities, I prove exact relations $N_{i}=2T-V+2$ and $T-N_{i}+S_{i}=1$ (with $S_{i}$ taken to be the number of interior edges) and obtain restricted linear ranges for internally decomposed polyhedra with the minimal number of added internal edges. Consequently, I propose a symbolic workflow that yields rapid pre-checks for structural impossibility, reducing the need for costly geometric validation in computational geometry, graphics, and automated modeling.
@@ 1 Introduction and Background
Scope and fixed assumptions: Throughout this work, I limit attention to genus-zero (simply connected) surfaces, with no geometric embeddability tests (i.e., no attempt to prove realizability in $\mathbb{R}^{3}$), and without consideration of higher-genus or non-manifold topologies. Unless stated otherwise, all internal “range” results are given under the strict internal decomposition procedure described in the Methods section, i.e. normal form with at most one new interior segment per layer, introduced only when necessary. All identities and bounds presented are symbolic rather than metric, and some are conjectural or supported by empirical validation rather than proved in full generality.
@@
Combining these with the constraint that all vertices have degree at least $3$ yields sharp bounds:
@@
The auxiliary variables $N$, $M$, $S$, $N_{i}$, and $S_{i}$ do not appear in the classical Euler–Steinitz framework. They are introduced here to support symbolic decomposition bookkeeping and to connect surface composition to volumetric tetrahedralization. N represents the total internal angle units on the external surface, with one unit defined as $\pi$ radians ($180^{\circ}$ ). $M$ is derived from polygon types and measures “external angle fullness” in the same units. $S$ counts the non-intersecting triangulation segments added to polygonal faces to produce a fully triangulated surface and also serves as a discrete flatness measure. $N_{i}$ counts the number of internal gluing angle units, each corresponding to a shared triangular face between two tetrahedra. $S_{i}$ counts the internal non-intersecting triangulation segments required to complete a volumetric mesh during decomposition. None of these auxiliary quantities are edges in the polyhedron’s graph; they exist only in the symbolic model and may not correspond to realizable geometric features.
@@ 1.1 Illustrative Examples of $N_{i}$, $S$, and $S_{i}$
Figure 1: Rectangular pyramid illustrating the flatness measure $S$. The single internal diagonal on the quadrilateral base (orange) is the unique surface triangulation segment needed to obtain a fully triangulated boundary, hence $S=1$. This corresponds to one coplanarity merge relative to the maximally triangulated case and, by Proposition 4.3, decreases both $F$ and $E$ by $1$ while preserving $V$ and $\chi=2$.
@@
2 Preliminaries
2.1 Normal form and standing assumptions
Definition 2.1 (Normal form).
A genus-0 polyhedron is in normal form if: (i) all vertices lie on the boundary (no interior/Steiner vertices, so $V_{i}\equiv 0$); (ii) edges are straight segments that intersect only at shared endpoints (boundary $1$-skeleton is a planar straight-line graph); (iii) every interior edge is an internal triangulation segment between boundary vertices and is incident to exactly two distinct triangular faces; and (iv) internal adjacencies occur only along triangular faces (no edge bisections).
Remark (Notation and scope).
Throughout, $S:=\sum_{f}(\deg f-3)$ denotes flatness. Unless stated otherwise, all identities tagged “Proposition” are proved under normal form; anything tagged “Heuristic” or “Conjectural Proposition” is indicated where used.
@@ 3.6 Tetrahedral decomposition and trivalent vertex prioritization
This work restricts attention to genus–0, normal–form decompositions built by peeling one outer shell of tetrahedra at a time, a procedure I will refer to informally as a shell–aligned ladder tetrahedralization (SALT). Each layer is created by coning existing boundary triangles and, only if necessary, introducing at most one new interior segment between boundary vertices to complete gluing across that layer. No interior vertices (Steiner points) and no edge bisections are allowed, and internal adjacencies occur only along triangular faces.
Within this scheme, a “minimal interior edge” (MIE) rule is observed: a new interior segment is introduced only when strictly necessary to complete the current layer; otherwise, none is added.
@@ 4.1 Vertex Basis and Initial Angle Unit Calculation
For a genus-0 polyhedron in normal form, consider first the triangulated case ($S=0$). In this case every face has degree $3$, so the incidence relation is
@@
Now, increasing the flatness parameter $S$ by one corresponds to replacing a triangular face by a face of degree $4$; this removes exactly one edge and one face from the triangulation. More generally, each unit of $S$ reduces both $E$ and $F$ by exactly one compared to the triangulated case. Therefore
@@
Let $G$ be the $1$-skeleton of a genus–0 polyhedron (simple, planar, and $3$-connected) with $V\geq 4$ vertices and $E$ edges. Then
@@
    $V$ is even and all vertices are trivalent (degree $3$); or
@@
    $V$ is odd and all but one vertex are trivalent, the remaining vertex being $4$-valent.
@@
Every vertex has degree at least $3$, so by the handshake lemma
@@
If $V$ is even and all vertices have degree $3$, then $2E=3V$ and $E=3V/2$, attaining the bound.
If $V$ is odd and all but one vertex are trivalent and the remaining vertex is $4$-valent, then 2E=3(V1)+4=3V+12E=3(V-1)+4=3V+1, so $E=(3V+1)/2$, again attaining the bound.
@@
Moreover, as $S$ increases by $1$ (flattening step), both $E$ and $F$ decrease by $1$, and for $k=0,1,\dots,S_{\max}$ one has E=3V6k,F=2V4k.E=3V-6-k,\;F=2V-4-k.
@@
Start from any maximally planar (triangulated) 33–connected planar graph on $V$ vertices (e.g., a stacked triangulation), which has $S=0$, E=3V6E=3V-6, F=2V4F=2V-4. A “coplanarity merge” of two adjacent boundary triangles corresponds combinatorially to deleting their shared edge, which decreases both $E$ and $F$ by $1$ and keeps $V$ fixed. In a stacked triangulation one can choose a set of V4V-4 edges (e.g., along successive stacked faces) whose deletion preserves simplicity, planarity, and 33–connectivity; performing $S$ such merges yields a graph with the claimed $(E,F)$. ∎
@@ 4.3 Configuration Series over Flatness
From Proposition 4.1, each unit decrease in the flatness measure $S$ increases both $F$ and $E$ by $1$. Starting from the extremal values in Corollary 4.2, I have:
@@ 4.4 Counting Distinct Combinatorial Configurations
Heuristic 4.1 (External bounds and parity effects).
@@
since each tetrahedron contributes four triangle units and every internal gluing triangle removes two boundary triangle units by pairing. These two expressions for $N$ are consistent with Proposition 6.4 and Proposition 4.1: indeed $F+S=4T-2N_{i}$. This connection between external angle counts and internal tetrahedral composition will be used in §5 to link internal volumetric parameters directly to external face and edge counts.
@@ 4.5 Triangulation Sequences and the Invariance of Euler’s Characteristic
If $S$ denotes the number of coplanarity merges of adjacent triangles (each merge decreases both $F$ and $E$ by $1$ while $V$ is fixed), then for all $S$ in the admissible range,
@@
Each triangle has three edges and each edge is shared by two triangles, so $2E_{\max}=3F_{\max}$. Substituting $E_{\max}=\tfrac{3}{2}F_{\max}$ into Euler’s formula $V-E_{\max}+F_{\max}=2$ yields $F_{\max}=2(V-2)$ and $E_{\max}=3(V-2)$. A coplanarity merge of two adjacent triangles removes one surface edge and one triangle, so $F\mapsto F-1$ and $E\mapsto E-1$ with $V$ unchanged. Hence VE+FV-E+F is invariant and equal to $2$ for all admissible $S$. ∎
@@ 5 Internal Decomposition and Tetrahedral Estimation
Scope disclaimer.
Throughout this section, “SALT+MIE” denotes a restricted construction class (genus–0, normal form; shell layers; no edge bisections). Statements labeled as identities are exact under the stated hypotheses; ranges and patterns specific to SALT+MIE are not claimed to hold for arbitrary tetrahedralizations.
@@ 5.4 Count of Valid Internal Configurations
Heuristic 5.1 (Internal ranges and configuration ladder).
@@ 6.1 Exact topological Euler link
  • Faces: $N+N_{i}$, where $N$ is the number of boundary triangular faces (angle units),
@@ 6.3 Heuristic extension of Euler’s characteristic
Heuristic 6.1 (Extended Euler-type identity).
@@ 7 A Combinatorial Method for Testing 3D Polygonal Enclosure and Polyhedral Feasibility
If $N$ is odd, then $V$ is non-integer and no valid genus-0 polyhedron in normal form with such a face–degree distribution can exist.
@@ 11 Acknowledgements
This appendix records (i) proved external and topological identities used in the main text, and (ii) the empirical/heuristic identities and bounds collected from worked decompositions. All variables are symbolic and coordinate-free. Genus 0 and normal-form assumptions apply where stated. All linear “max” values quoted below are valid within the SALT+MIE class. Non–SALT triangulations can exceed these bounds.
@@ E.1 Proved External Identities and Bounds
(Equation (E.6) counts triangle units: each tetrahedron contributes $4$, each internal gluing triangle cancels $2$.)
@@ E.2 Proved Topological Identity for 3D Complexes
(Prop. 6.1). In particular, when $V_{i}=0$ (normal form, no interior vertices),